Colombo City typically has rain on about 16% of days in November.
(a) What is the probability that it will rain on exactly 5 days in November? 15 days?
(b) What is the mean number of days with rain in November?
(c) What is the variance and standard deviation of the number of days with rain in November?
1
Expert's answer
2019-12-26T10:07:37-0500

Probability of having the rain(p)=16%=0.16

Probability of not having the rain(q)=1-p=1-0.16=0.84

No of days in the month of November =30 days

(a) i) The probability that it will rain on exactly 5 days in November

=nCrprqn-r=30C5(0.16)5(0.84)30-5=

=142,506 x 0.0001048576 x 0.0127932873

=0.182

ii) The probability that it will rain on exactly 15 days in November=

30C15(0.16)15(0.84)30-15=

=155,117,520 x 8.43313454 x 10-14

=1.30812692 x 10-5

b) mean number =np=30x0.16=4.8

c) The variance and standard deviation of the number of days with rain in November

variance= npq= 30 x 0.16 x0.84=4.032


standard deviation = "\sqrt{variance}=\sqrt{4.032}=2.00798406."